# A new perspective on modulus stabilization

### v2

@@ Abstract
In the original Randall-Sundrum (RS) two-brane model, the radion potential was fine-tuned to zero, leaving the extra-dimensional modulus undetermined. Goldberger and Wise (GW) introduced a stabilization mechanism by incorporating a massive bulk scalar field, leading to a radion potential with a stable minimum. In this work, we examine the impact of the intrinsic stability of the GW scalar field on the modulus stabilization procedure. To elucidate this relationship, we test the stabilization scheme across several relevant scenarios using methods of singular perturbation theory. Our analysis consistently indicates a one-to-one correspondence between the stability of the GW scalar and successful modulus stabilization. Additionally, we highlight that certain classes of GW bulk potentials demonstrate such a correspondence explicitly, and provide checks which any GW bulk potential must satisfy in order to ensure modulus stabilization.
preprint: APS/123-QED
I Introduction
@@
The gauge hierarchy problem in Standard Model (SM) of particle physics results into the well known fine-tuning problem in connection to the Higgs mass which acquires a quadratic divergence due to the large radiative corrections in perturbation theory. In order to confine the Higgs mass parameter within TeV scale, one needs to consider theories beyond SM. Among many such attempts, the two-brane Randall-Sundrum (RS) warped extra-dimensional scenario has earned special attention for following reasons [1]: (a) it potentially resolves the gauge hierarchy problem without introducing any other intermediate scale in the theory and (b) the extra-dimensional modulus can be stabilized by introducing a bulk scalar field in the setup [2].
@@
One of the crucial aspects of this braneworld model is to stabilize the distance between the two branes (known as modulus or radion). For this, one needs to generate an appropriate potential for the radion field with a stable minimum consistent with the value proposed in RS model in order to solve the gauge hierarchy problem. Goldberger and Wise (GW) [2, 3] proposed a very useful mechanism to achieve this by introducing a bulk scalar field with suitable bulk and brane potentials. They showed that one can indeed stabilize the modulus without any unnatural fine-tuning of the parameters though they ignored effects of the backreaction of the bulk scalar on the background metric.
@@
Since GW, different stabilization procedures have been proposed and among these, DeWolfe et al. [4] is one of the most important ones as it provided an exact solve including the backreaction, which was missing in the original GW proposal. Later on, Chacko et al. [5, 6] revisited the GW stabilization including the gravitational contribution to radion potential and also including self interaction terms for the bulk potential, though neglecting backreaction. They heavily relied on an approximate solution for the scalar field profile (for a non-trivial bulk potential) and this is something we have discussed in some detail here as well. We also highlight that several variants of the RS model, their respective modulus stabilization and cosmological implications were explored previously in Refs. [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Quantum effects in such theory were studied in Refs. [19, 20].
@@
Given all this research, an evident question to ask is how does the scalar field profile - in other words, the stability of the scalar field affect the modulus stabilization procedure? Is there a one-one correspondence between the two stabilities? This is the problem we set out to explore in this work. Apparently, the relationship between the two stabilities seems highly non-trivial because of the sheer number of steps involved in reaching upto the radion potential starting from the scalar field profile. So, the way we wish to tackle this problem is by exploring the modulus stabilization scheme taking different forms for the scalar action, incorporating different forms for the bulk potentials.
The primary scenarios that we discuss and provide a detailed analysis for are the following:
  • Flipped signs in the original GW proposal i.e. a phantom or tachyon instability.
  • Bulk cubic interaction.
  • Bulk quartic interaction.
  • Bulk double-well potential.
  • Bazeia-Furtado-Gomes (BFG) type bulk potentials.
  • Mishra-Randall ansatz bulk potentials.
Each of these cases is discussed in depth and most of them (as we will see eventually) hint at the one-one correspondence we mentioned. We also prove the heightened possibility of such a correspondence in some classes of bulk scalar potentials that further solidify our intuition. Lastly, we provide a simplified criterion that any general GW bulk potential must satisfy to enable modulus stabilization.
@@
The structure of the paper is as follows. Sec.II provides the RS-I setup and lays out the original GW prescription. Sec.III discusses a more general form of the GW scheme involving gravitational contributions to radion potential and bulk interaction terms for the GW scalar. Sec.IV deals with the six different scenarios we mentioned previously  [^1] It is to be noted that when we talk of the one-one correspondence between radion stabilization and GW scalar stability, we always take the radion potential generated only due to the GW part and not the garvitational sector contribution, unless mentioned otherwise.. Sec.V presents the GW potential classes where this correspondence is highly probable. Sec.VI provides the criterion on any bulk scalar potential to achieve modulus stabilization. A short conclusion is provided in Sec.VII.
II RS-I and GW
II.1 The RS-I Model
@@
    $\begin{split}ds^{2}=G_{MN}dx^{M}dx^{N}=e^{-2\sigma}\eta_{\mu\nu}dx^{\mu}dx^{\nu}\\ -r_{c}^{2}d\phi^{2}\end{split}$ (1)
    where $\sigma$ is the warp factor, the greek index runs from 0 to 3 while the latin index runs from 0 to 5, excluding 4 with 5 signifying the fifth spatial dimension, and $r_{c}$ is the radius of compactification of the extra dimension. We work with the mostly negative metric signature.
  • The topology of the extra dimension is $S^{1}/\mathbb{Z}_{2}$. The fifth coordinate is angular in nature with $(x,\phi)$ identified with $(x,-\phi)$.
  • The 3-branes extending in $x^{\mu}$ direction are located at $\phi=0$ and $\phi=\pi$
  • The induced metrics on the branes are given by:
@@
II.2 Why modulus stabilization?
The 5D RS action without any bulk scalar is given by Eq.(II.1). Starting from this 5D action involving gravity only, the effective 4D action involving the 4D graviton and radion fields is found out by plugging in the 5D Ricci scalar $\mathcal{R}$ and integrating over the extra dimension. Consequently, the RS metric that we use is (we promote the metric and modulus to be dynamical fields)
@@ II.2 Why modulus stabilization?
In absence of the 4D graviton, $g_{\mu\nu}\rightarrow\eta_{\mu\nu}$ as in the original RS prescription. Defining $\varphi\equiv Ae^{-k\pi r(x)}$, where $A=\sqrt{24M^{3}/k}$ to be the canonical radion field, one obtains the following form for the gravitational contribution to radion potential [5]
@@
where $\Lambda_{4D}=\Big{(}T_{hid}+\frac{\Lambda_{b}}{k}\Big{)}/k^{4}$ is the $4D$ cosmological constant which we tune to be small, and $\tau=\Big{(}T_{vis}-\frac{\Lambda_{b}}{k}\Big{)}/k^{4}$. Thus, the potential is essentially $V_{GR}(\varphi)=\tau(k\varphi/A)^{4}$. Originally, [1] set this potential to zero through fine tunings of the brane tensions i.e. $T_{vis}=-T_{hid}=\Lambda_{b}/k$ (see Eq.6). However, then the radion field could take up any possible value. To fix this, one needed a stabilization mechanism that fixes the interbrane separation and makes the radion acquire a mass.
II.3 Goldberger-Wise Stabilization
In a realistic RS scenario, a mechanism is needed to stabilize the geometry and give the radion a mass. Such a mechanism was proposed by Goldberger and Wise [2]. In this construction, a massive $5D$ field $\Phi$ is sourced at the boundaries and acquires a VEV whose value depends on the location in the extra dimension. After integrating over the extra dimension, this generates a radion potential in the low energy effective theory. In the original GW construction, the quartic potential for the radion from the gravity sector was tuned to zero, and only the dynamics of the scalar field $\Phi$ contributed to the radion potential.
@@ II.3 Goldberger-Wise Stabilization
[…]
@@
We get a $\phi$-dependent vacuum expectation value $\Phi(\phi)$ which is determined classically by solving the differential equation (Euler-Lagrange equation for the scalar field $\Phi$),
$\begin{split}-{1\over r_{c}^{2}}\partial_{\phi}(e^{-4\sigma}\partial_{\phi}\Phi)+m^{2}e^{-4\sigma}\Phi+4e^{-4\sigma}\lambda_{v}\Phi(\Phi^{2}-v_{v}^{2})\\ \frac{\delta(\phi-\pi)}{r_{c}}\mbox{}+4e^{-4\sigma}\lambda_{h}\Phi(\Phi^{2}-v_{h}^{2})\frac{\delta(\phi)}{r_{c}}=0\end{split}$ (13)
where $\sigma(\phi)=kr_{c}|\phi|$. The bulk solution for $\Phi$ is then given by
@@
with $\nu=\sqrt{4+m^{2}/k^{2}}$. Plugging this solution back into the scalar field action and integrating over $\phi$ yields an effective four-dimensional potential $V_{\Phi}$ for $r_{c}$ which involves the constants A and B. These are determined from the boundary conditions i.e. matching with the delta functions. In the limit of large $\lambda_{h}$, $\lambda_{v}$ and $kr_{c}$, by taking $\Phi(0)=v_{h}$ and $\Phi(\pi)=v_{v}$, we get
$A=v_{v}e^{-(2+\nu)kr_{c}\pi}-v_{h}e^{-2\nu kr_{c}\pi}$ (15)
$B=v_{h}(1+e^{-2\nu kr_{c}\pi})-v_{v}e^{-(2+\nu)kr_{c}\pi}$ (16)
where subleading powers of $\exp(-kr_{c}\pi)$ have been neglected. Suppose that $m/k\ll 1$ so that $\nu=2+\epsilon,$ with $\epsilon\approx m^{2}/4k^{2}$ a small quantity. Plugging these back into the expression for $V_{\Phi}(r_{c})$ and taking the large $kr_{c}$ limit, the potential becomes
$\begin{split}V_{\Phi}(r_{c})=k\epsilon v_{h}^{2}+4ke^{-4kr_{c}\pi}(v_{v}-v_{h}e^{-\epsilon kr_{c}\pi})^{2}\left(1+\frac{\epsilon}{4}\right)\\ -k\epsilon v_{h}e^{-(4+\epsilon)kr_{c}\pi}(2v_{v}-v_{h}e^{-\epsilon kr_{c}\pi})\end{split}$ (17)
where terms of order $\epsilon^{2}$ are neglected though $\epsilon kr_{c}$ is not neglected. Ignoring terms proportional to $\epsilon$, this potential has a minimum at
@@
III Generalized Goldberger-Wise: CMS Method
In the original Goldberger-Wise (GW) construction, the quartic potential for the radion $V_{GR}(\varphi)$ was fine-tuned to zero, leaving the dynamics of the scalar field $\Phi$ as the sole contributor to the radion potential. However, it is possible to relax this constraint by allowing a non-zero quartic contribution from the gravitational potential and incorporating self-interaction terms in the bulk potential of the GW scalar. This generalized approach has been effectively explored by Chacko, Mishra, and Stolarski in Ref.[5]. We henceforth call this approach the CMS method, after the authors of Ref.[5].
III.1 Approximate solution for the scalar field profile
@@
$\begin{split}\mathcal{S}_{GW}=\int d^{4}x\>d\phi\>[\sqrt{G}\>(\frac{1}{2}G^{AB}\partial_{A}\Phi\partial_{B}\Phi-V_{b}(\Phi))\\ -\sum_{i=\{vis,hid\}}\delta(\phi-\phi_{i})\>\sqrt{-G_{i}}\>V_{i}(\Phi)]\;\end{split}$ (19)
@@
where $\alpha$ is a dimensionless number. As Ref.[5] states, the choice of a linear potential for $\Phi$ on the visible brane is natural and is generally expected when $\Phi$ is not charged under any symmetries. The bulk potential for the GW scalar $\Phi$ has the general form
@@
$\phi=\pi\qquad:\qquad\partial_{\phi}\Phi=-\alpha k^{3/2}kr_{c}\;\;$ (22)
We consider solutions for $0\leq\phi\leq\pi$ because of the ${\mathbb{Z}}_{2}$ orbifold symmetry. It follows that in the limit of large $4kr_{c}$, two independent approximate solutions to Eq.(III.1) can be obtained in the following way: once by dropping the $\partial_{\phi}^{2}\Phi$ term, and once by dropping the potential term. This is motivated by methods of singular perturbation theory. These two equations are denoted as the outer region (OR) and boundary region (BR) equations respectively. The OR solution holds in the bulk, while the BR solution holds close to the visible brane where a boundary layer is formed with thickness of the order $\sim 1/4kr_{c}$.
[…]
The BR solution being independent of the bulk potential $V_{b}$ is readily solved. On applying the boundary condition at $\phi=\pi$, the BR solution is given by
@@
The constant $C$ is determined by asymptotic matching to the OR solution [21]. Hence, we get a smooth solution that very well approximates the two independent solutions in their regions of relevance. The parameters $\alpha$ and $v$ are chosen to be small and almost of the same order.
III.2 Getting the stabilized modulus
On getting the approximate scalar field solution, we plug it back into the GW action (Eq.19) along with the corresponding bulk and brane potentials. We then integrate over the extra dimension, add gravitational contributions to it i.e. $V_{GR}(\varphi)$ and work up to the leading order to get the radion potential. Effectively, in presence of the GW scalar, the $\tau$ in $V_{GR}(\varphi)$ will receive additional corrections. At the minimum of this radion potential, we say that the modulus is stabilized and the corresponding value of $r_{c}$ is what we seek (see Appendix B, [5] for details).
IV In search of a stabilized modulus
IV.1 Flipping signs in the original GW setup
Suppose we flip the sign of the kinetic term (phantom instability) or the mass term (tachyon instability) in Eq. (10). We wish to now know whether we still find a stabilized modulus for our model. We should follow same steps as previously and write down the new classical equation of motion for the bulk scalar,

### v4

# Analyzing the general conditions for modulus stabilization in a warped braneworld

@@ Abstract
[…]
1 Introduction
@@
The gauge hierarchy problem in Standard Model (SM) of particle physics results into the well known fine-tuning problem in connection to the Higgs mass which acquires a quadratic divergence due to the large radiative corrections in perturbation theory. In order to confine the Higgs mass parameter within TeV scale, one needs to consider theories beyond SM. Among many such attempts, the two-brane Randall-Sundrum (RS) warped extra-dimensional scenario has earned special attention for following reasons RS1: (a) it potentially resolves the gauge hierarchy problem without introducing any other intermediate scale in the theory and (b) the extra-dimensional modulus can be stabilized by introducing a bulk scalar field in the setup GW1.
@@
One of the crucial aspects of such braneworld models is to stabilize the distance between the two branes (known as modulus or radion). For this, one needs to generate an appropriate potential for the radion field with a stable minimum consistent with the value proposed in RS model in order to solve the gauge hierarchy problem. Goldberger and Wise (GW) GW1; GW2 proposed a very useful mechanism to achieve this by introducing a bulk scalar field with suitable bulk and brane potentials. They showed that one can indeed stabilize the modulus without any unnatural fine-tuning of the parameters, though ignoring effects of the backreaction of the bulk scalar on the background metric.
@@
Since the GW proposal, different stabilization procedures have been proposed and among these, DeWolfe et al. Freedman is one of the most prominent ones as it provided an exact solve including the backreaction, which was missing in the original GW prescription. Later on, Chacko et al. Chacko1; Chacko2 revisited the GW stabilization including the gravitational contribution to radion potential and also including self interaction terms for the bulk potential, albeit neglecting backreaction. They heavily relied on an approximate solution for the scalar field profile (for a non-trivial bulk potential) and this is something we have discussed in some detail here as well. Several variants of the RS model, their respective modulus stabilization and cosmological implications were explored previously in Refs. csaki2001; gwvariations; geometry_modulus; SauryaDas_2008; PhysRevD.75.107901; Das_mukherjee; tanmoy2; banerjeepaul; Banerjeessg; tanmoy; inflation_ssg; Randall_2023; ashmitassg; Koley_2009. Quantum effects in such theory were studied in Refs. odinstov1; odinstov2.
@@
Given all this research, an evident question to ask is how does the scalar field profile affect the modulus stabilization procedure? Is there a correspondence between the nature of the GW bulk potential and successful modulus stabilization? This is the problem we set out to explore in this work. Apparently, the relationship between the two seems highly non-trivial because of the sheer number of steps involved in reaching upto the radion potential starting from the scalar field profile. So, the way we wish to tackle this problem is by exploring the modulus stabilization scheme taking different scalar actions, incorporating different forms for the bulk potentials. It must be noted that in AdS space which is the nature of the bulk in Randall-Sundrum, the stability of the GW field only requires satisfying the Breitenlohner-Freedman (BF) bound 1982: $m^{2}/k^{2}+4>0$ where $m^{2}$ is the mass squared of the GW field and $k$ is the AdS curvature scale. Thus, though the presence of a minimum in the GW bulk potential clearly indicates a stable field; the lack of a minimum does not mean an unstable GW field as long as the BF bound is satisfied.
The primary scenarios that we discuss and provide a detailed analysis for are the following: (a) stabilizing scalar with no bulk interactions, (b) bulk cubic interaction, (c) bulk quartic interaction, (d) bulk double-well potential, (e) Bazeia-Furtado-Gomes (BFG) type bulk potentials, and (e) Mishra-Randall ansatz bulk potentials. We highlight a correspondence between the nature of the bulk scalar potential and modulus stabilization. Lastly, we provide a simplified criterion that any bulk scalar potential must satisfy to enable modulus stabilization.
The other aspect of the paper is the following: it is a known fact that checking for modulus stabilization and thereby getting to the stabilized modulus value is cumbersome for any non-trivial bulk potential - we solve for the scalar field profile - plug it back into the GW action - integrate over the extra dimension - add gravitational contributions (if necessary) to it - and work up to the leading order to get the radion potential. So, a natural question to ask is whether there exists an alternative approach that could yield the stabilized modulus value (if it gets stabilized), at least in certain cases? Here we show that one could indeed avoid all the hassle and still get a bound on the stabilized modulus ($r_{c}$), employing the brane world sum rules, put down by Gibbons et al. in Ref. Gibbons in certain scenarios. We also show that these bounds are actually close to the true stabilized modulus values that one gets by minimizing the radion potential. This is especially interesting as it might serve as a go-to prescription for estimating the stabilized modulus value and a hint at possibly unexplored physics.
[…]
@@
The structure of the paper is as follows. Sec.2 provides a quick review of the RS-I setup and the original GW prescription. Sec.3 discusses a more general form of the GW scheme involving gravitational contributions to radion potential and bulk interaction terms for the GW scalar. Sec.4 deals with the six different scenarios we mentioned previously  [^1] It is to be noted that when we talk of radion stabilization, we always take the radion potential generated only due to the GW part and not the gravitational sector contribution, unless mentioned otherwise (Sec.7 includes the gravitational contribution).. Sec.5 highlights the correspondence between the nature of bulk scalar potentials and successful radion stabilization. Sec.6 provides the criterion on any bulk scalar potential to achieve modulus stabilization. Sec.7 discusses the connection with brane world sum rules. A short conclusion is provided in Sec.8.
2 Warped extra-dimensions and modulus stabilization
2.1 Randall-Sundrum I
@@
$\begin{split}ds^{2}=G_{MN}dx^{M}dx^{N}=e^{-2\sigma}\eta_{\mu\nu}dx^{\mu}dx^{\nu}-r_{c}^{2}d\phi^{2}\end{split}$ (1)
where $\sigma$ is the warp factor, the greek index runs from 0 to 3 while the latin index runs from 0 to 5, excluding 4 with 5 signifying the fifth spatial dimension, and $r_{c}$ is the radius of compactification of the extra dimension. We work with the mostly negative metric signature. The topology of the extra dimension is $S^{1}/\mathbb{Z}_{2}$. The fifth coordinate is angular in nature with $(x,\phi)$ identified with $(x,-\phi)$. The 3-branes extending in $x^{\mu}$ direction are located at $\phi=0$ and $\phi=\pi$. The induced metrics on the branes are given by:
@@
2.2 Why modulus stabilization?
The 5D RS action without any bulk scalar is given by Eq.(2.1). Starting from this 5D action involving gravity only, the effective 4D action involving the 4D graviton and radion fields is found out by plugging in the 5D Ricci scalar $\mathcal{R}$ and integrating over the extra dimension. Consequently, the RS metric that we use is (we promote the metric and modulus to be dynamical fields)
@@ 2.2 Why modulus stabilization?
In absence of the 4D graviton, $g_{\mu\nu}\rightarrow\eta_{\mu\nu}$ as in the original RS prescription. Defining $\varphi\equiv Ae^{-k\pi r(x)}$, where $A=\sqrt{24M^{3}/k}$ to be the canonical radion field, one obtains the following form for the gravitational contribution to radion potential Chacko1
@@
where $\Lambda_{4D}=\Big(T_{hid}+\frac{\Lambda_{b}}{k}\Big)/k^{4}$ is the $4D$ cosmological constant which we tune to be small, and $\tau=\Big(T_{vis}-\frac{\Lambda_{b}}{k}\Big)/k^{4}$. Thus, the potential is essentially $V_{GR}(\varphi)=\tau(k\varphi/A)^{4}$. Originally, RS1 set this potential to zero through fine tunings of the brane tensions i.e. $T_{vis}=-T_{hid}=\Lambda_{b}/k$ (see Eq.6). However, then the radion field could take up any possible value. To fix this, one needed a stabilization mechanism that fixes the interbrane separation and makes the radion acquire a mass. This is essential as the stabilized modulus enters as a parameter in the lower dimensional effective theory.
2.3 Goldberger-Wise stabilization
In a realistic RS scenario, a mechanism is needed to stabilize the geometry and give the radion a mass. Such a mechanism was proposed by Goldberger and Wise GW1. In this construction, a massive $5D$ field $\Phi$ is sourced at the boundaries and acquires a VEV whose value depends on the location in the extra dimension. After integrating over the extra dimension, this generates a radion potential in the low energy effective theory. In the original GW construction, the quartic potential for the radion from the gravity sector was tuned to zero, and only the dynamics of the scalar field $\Phi$ contributed to the radion potential.
@@ 2.3 Goldberger-Wise stabilization
$S_{b}={1\over 2}\int d^{4}x\int_{-\pi}^{\pi}d\phi\sqrt{G}\left(G^{AB}\partial_{A}\Phi\partial_{B}\Phi-m^{2}\Phi^{2}\right)$ (10)
@@
We get a $\phi$-dependent vacuum expectation value $\Phi(\phi)$ which is determined classically by
@@
where $\sigma(\phi)=kr_{c}|\phi|$ with $\nu=\sqrt{4+m^{2}/k^{2}}$. Plugging this solution back into the scalar field action and integrating over $\phi$ yields an effective four-dimensional potential $V_{\Phi}$ for $r_{c}$ which involves the constants A and B. These are determined from the boundary conditions i.e. matching with the delta functions. In the limit of large $\lambda_{h}$, $\lambda_{v}$ and $kr_{c}$, $\Phi(0)=v_{h}$ and $\Phi(\pi)=v_{v}$ are the energetically favoured boundary conditions. In the large $kr_{c}$ limit, the radion potential turns out to be
$\begin{split}V_{GW}(r_{c})=k\epsilon v_{h}^{2}+4ke^{-4kr_{c}\pi}(v_{v}-v_{h}e^{-\epsilon kr_{c}\pi})^{2}\left(1+\frac{\epsilon}{4}\right)-k\epsilon v_{h}e^{-(4+\epsilon)kr_{c}\pi}(2v_{v}-v_{h}e^{-\epsilon kr_{c}\pi})\end{split}$ (14)
where $m/k\ll 1$ so that $\nu=2+\epsilon,$ with $\epsilon\approx m^{2}/4k^{2}$ a small quantity. Terms of order $\epsilon^{2}$ are neglected though $\epsilon kr_{c}$ is not neglected. Ignoring terms proportional to $\epsilon$, this potential has a minimum at
@@
3 Generalized Goldberger-Wise stabilization
In the original Goldberger-Wise (GW) construction, the quartic potential for the radion $V_{GR}(\varphi)$ was fine-tuned to zero, leaving the dynamics of the scalar field $\Phi$ as the sole contributor to the radion potential. However, it is possible to relax this constraint by allowing a non-zero quartic contribution from the gravitational potential and incorporating self-interaction terms in the bulk potential of the GW scalar. This generalized approach has been effectively explored by Chacko, Mishra, and Stolarski in Ref.Chacko1. We henceforth call this approach the CMS method, after the authors of Ref.Chacko1.
3.1 Approximate solution for the scalar field profile
@@
$\begin{split}\mathcal{S}_{GW}=\int d^{4}x\>d\phi\>[\sqrt{G}\>(\frac{1}{2}G^{AB}\partial_{A}\Phi\partial_{B}\Phi-V_{b}(\Phi))-\sum_{i=\{vis,hid\}}\delta(\phi-\phi_{i})\>\sqrt{-G_{i}}\>V_{i}(\Phi)]\;\end{split}$ (16)
@@
where $\alpha$ is a dimensionless number. As is known and Ref.Chacko1 states as well, the choice of a linear potential for $\Phi$ on the visible brane is natural and is generally expected when $\Phi$ is not charged under any symmetries. For example, if $\Phi$ were charged under a $\mathbb{Z}_{2}$ symmetry i.e. $\Phi\rightarrow-\Phi$, then the linear term would be forbidden, and the leading allowed term would be quadratic or higher.
The bulk potential for the GW scalar $\Phi$ has the general form
@@
$\phi=\pi\qquad:\qquad\partial_{\phi}\Phi=-\alpha k^{3/2}kr_{c}\;\;\$(19)
We consider solutions for $0\leq\phi\leq\pi$ because of the ${\mathbb{Z}}_{2}$ orbifold symmetry. It follows that in the limit of large $4kr_{c}$, two independent approximate solutions to Eq.(3.1) can be obtained in the following way: once by dropping the $\partial_{\phi}^{2}\Phi$ term, and once by dropping the potential term. This is motivated by methods of singular perturbation theory. These two equations are denoted as the outer region (OR) and boundary region (BR) equations respectively. The OR solution holds in the bulk, while the BR solution holds close to the visible brane where a boundary layer is formed with thickness of the order $\sim 1/4kr_{c}$.
$\frac{d\Phi}{d\phi}=-\frac{r_{c}}{4k}V_{b}^{\prime}(\Phi)\;\;\;\;\;\;\;\;\textbf{(OR)}$
$\frac{d^{2}\Phi}{d\phi^{2}}=4kr_{c}\frac{d\Phi}{d\phi}\;\;\;\;\;\;\;\;\textbf{(BR)}$ (20)
The BR solution being independent of the bulk potential $V_{b}$ is readily solved. On applying the boundary condition (b.c.) at $\phi=\pi$, the BR solution is given by
@@
The constant $C$ is determined by asymptotic matching to the OR solution Bender2010-wx. Hence, we get a smooth solution that very well approximates the two independent solutions in their regions of relevance. The parameters $\alpha$ and $v$ are chosen to be small and almost of the same order.
3.2 Getting the stabilized modulus
On getting the approximate scalar field solution, we plug it back into the GW action (Eq.16) along with the corresponding bulk and brane potentials. We then perform dimensional reduction, add gravitational contributions to it i.e. $V_{GR}(\varphi)$ and work up to the leading order to get the radion potential. Effectively, in presence of the GW scalar, the $\tau$ in $V_{GR}(\varphi)$ will receive additional corrections. At the minimum of this radion potential, we say that the modulus is stabilized and the corresponding value of $r_{c}$ is what we seek (see Appendix B, Chacko1 for details).
4 In search of a stabilized modulus
4.1 No bulk interactions
In the Goldberger-Wise prescription, they used a stabilizing scalar with no bulk interaction. But now suppose that we flip the sign of the kinetic term (phantom instability) or the mass term (tachyon instability) in Eq. (10). We wish to now know whether we still find a stabilized modulus for our model. We should follow same steps as previously and write down the new classical equation of motion for the bulk scalar,